A body travels 200cm in the first 2sec and 220cm in the next 4sec with deceleration.the velocity of the body at the end of 7th second is

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,

let the initial velocity be u
deceleration be a
velocity after 2 sec=v1=u-at
=>v1=u-2a
u2-v2=2aS
=>u2-(u-2a)2=2xax200
=>a=u-s/2
=>a=u-100
Putting equations of motion for the next 4s
u2-v2=2xax220
=>(u-2a)2-v2=2(u-100)x220
=>v=√((200-u)^2-440(u-100))  (this is the final velocity after 6 seconds which can be the initial velocity for the 7th second)
Velocity after the 7th second=u-at
=√((200-u)^2-440(u-100)) – (u-100) m/s
this is the final velocity.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)